OpenStudy (anonymous):

You invest $3,000 in an account that has an annual interest rate of 2.5%, compounded continuously. How much money will be in the account after 5 years?

OpenStudy (anonymous):

Last one! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Now we return to the formula A = Pe^(rt) because we're compounding continuously again. In this case, P = 3000, r = 0.025 and t = 5, so this means... A = 3000*e^(0.025*5)

jimthompson5910 (jim_thompson5910):

do you remember what to do here?

OpenStudy (anonymous):

Could you please remind me what the value of e is?

jimthompson5910 (jim_thompson5910):

sure thing

jimthompson5910 (jim_thompson5910):

e is approximately 2.71828

jimthompson5910 (jim_thompson5910):

so A = 3000*e^(0.025*5) is approximately A = 3000*2.71828^(0.025*5)

OpenStudy (anonymous):

A= 3399

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

Fantastic! Thank you so so much for all your help again! =)

jimthompson5910 (jim_thompson5910):

anytime, glad I could help you out